# Second Generation 3-DAP 3-Dimensional Assignment Problem

### v2

@@ IV-B Power Adaptation of Existing Modulation
Given the signal model in (3), for compact presentation, we define $\bm{p}_{\ell}:=[p_{1,\ell},\ldots,p_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ and $\bm{h}_{\ell}:=[h_{1,\ell},\ldots,h_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ as the transmit power vector and channel vector at time slot $\ell$, respectively. Let $\tilde{\bm{x}}_{\ell}$ denotes the modulation vector containing a fixed modulation pattern, i.e., quadrature amplitude modulation (QAM). In the presence of fading, the received constellation point corresponding to the function output $f^{(i)}$ at time slot $\ell$ is then given by:
@@
$\tilde{\bm{v}}^{(i)}_{\ell}=\bm{a}_{i}^{\mathsf{T}}\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})\left(\bm{I}_{K}\otimes\mathds{1}_{Q}\right)\bm{p}_{\ell},\forall\ell\in[L],$
@@
where $\otimes$ denotes the Kronecker product. $\bm{I}_{Q}$ and $\bm{I}_{K}$ are identity matrices of size $Q\times Q$ and $K\times K$. $\mathds{1}_{Q}$ is a $Q\times 1$ vector with all elements equal to one, and the operator $\text{diag}(\cdot)$ constructs a diagonal matrix. To ensure valid computation, we incorporate the constraint in (6) and formulate the following optimization to minimize the power vectors $\bm{p}_{\ell}$:
$\mathcal{P}_{3}:=$ $\min_{\bm{p}_{\ell}}\quad\sum\nolimits_{\ell=1}^{L}\|\bm{p}_{\ell}\|_{2}^{2}$
@@
${\rm s.t.}\quad$ $\sum\nolimits_{\ell=1}^{L}|(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{B}_{\ell}\bm{p}_{\ell}|^{2}\geq\Delta f_{i,j},\quad\forall(i,j)\in[M]^{2},$
where $\bm{B}_{\ell}:=\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})(\bm{I}_{K}\otimes\mathds{1}_{Q})$. Similarly, by introducing the lifted matrix $\bm{P}_{\ell}:=\bm{p}_{\ell}\bm{p}_{\ell}^{\mathsf{H}}$ and dropping the rank constraint, a relaxed version of $\mathcal{P}_{3}$ can be formulated as:
$\mathcal{P}_{3}\approx\mathcal{P}_{4}:=$ $\min_{\bm{P}_{\ell}}\quad\sum\nolimits_{\ell=1}^{L}{\rm trace}(\bm{P}_{\ell})$
${\rm s.t.}\quad$ $\sum\nolimits_{\ell=1}^{L}{\rm Tr}(\bm{P}_{\ell}\bm{C}_{\ell}^{i,j})\geq\Delta f_{i,j},\quad\forall(i,j)\in[M]^{2},$
$\quad\bm{P}_{\ell}\succeq 0,\quad\forall\ell\in[L],$
where $\bm{C}_{\ell}^{i,j}:=\bm{B}_{\ell}^{\mathsf{H}}(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{B}_{\ell}$. After solving Problem $\mathcal{P}_{4}$ using a CVX solver [12], the power vector $\bm{p}^{*}_{\ell}$ can be recovered from the solution $\bm{P}^{*}_{\ell}$ either by applying Cholesky decomposition for an optimal solution, or by using Gaussian randomization for a suboptimal one [15]. These recovered power vectors are then used to adjust the magnitude and phase of the transmitted symbols accordingly.
@@ V Numerical Experiments
where $N_{s}$ is the number of Monte Carlo trials, $f^{(i)}$ is the desired function value, $\hat{f}^{(i)}_{j}$ is the estimated value in the $j$-th Monte Carlo trial. $f_{{\rm max}}$ and $f_{{\rm min}}$ denote the maximum and minimum values of the function output, respectively.
@@
Figure 2: Performance of SeMAC is evaluated across various SNRs in terms of NMSE with $K=8$ nodes.
V-A Performance of SeMAC
For the first experiment, we consider a network of $K=8$ nodes for computing the sum function $f=\sum_{k=1}^{K}x_{k}$, the product function $f=\prod_{k=1}^{K}x_{k}$ and the max function $f=\max_{k}x_{k}$, where $x_{k}\in\{1,2,3,4\}$. The results in Fig. 2 show that the NMSE decreases with increasing Signal-to-Noise Ratio (SNR), demonstrating the effectiveness of SeMAC. Increasing the number of time slots from $L=1$ to $L=2$ improves the computation accuracy by providing more degree of freedom for constellation design, though at the cost of higher latency.
@@ V-B Comparison to ChannelComp, ReMAC and Bit-Slicing
In this subsection, we compare ChannelComp with repetition, SeMAC, ReMAC, and Bit-Slicing for computing the product function with $K=4$ nodes over $L=2$ time slots, where input values are selected from $x_{k}\in\{1,2,\ldots,16\}$. As shown in Fig. 3, by exploiting diverse modulation patterns across time slots, SeMAC achieves lower NMSE by approximately 18 dB compared to ReMAC. While ChannelComp with repetition and ReMAC offer similar performance, ReMAC achieves greater energy efficiency by selectively repeating modulated symbols. Moreover, Bit-Slicing yields the highest NMSE, due to approximation errors introduced during the pre- and post-processing required to support nonlinear functions, as previously analyzed in [9].
@@ V-C Performance of SeMAC-PA
We analyze the performance of SeMAC-PA under Rayleigh fading with $K=8$ and QAM modulation of orders $Q=4$ and $Q=16$ for computing the product function over $L=2$ time slots. The fading coefficients $h_{k,\ell}$ are modeled as independent complex Gaussian random variables, i.e., $h_{k,\ell}\sim\mathcal{CN}(0,1)$. Fig. 4 shows that NMSE decreases with increasing SNR, demonstrating the effectiveness of power adaptation. However, SeMAC-PA relies on fixed QAM constellations, making it less effective than the optimized modulation design in SeMAC under channel inversion power control.
@@ VI Conclusion
In this letter, we proposed SeMAC, a multi-symbol modulation framework for digital AirComp that enables reliable function computation through distinct modulation patterns across multiple time slots. To support scenarios with fixed modulation, we introduced SeMAC-PA, which adapts transmit power and phase to reshape the received constellation. Numerical results showed that SeMAC outperforms ReMAC by up to 18 dB in nonlinear computation. Overall, the proposed SeMAC contributes toward the reliable computation in 6G systems.
-A Proof of Proposition 1
First, when $L\leq L^{*}_{k}$, assume that the matrix $\bm{X}_{k}^{*}$ is not the optimal solution, and there exists a matrix $\bm{\hat{X}}_{k}=\bm{X}^{*}_{k}+\bm{H}_{k}$ with $\mathrm{rank}(\bm{\hat{X}}_{k})\leq L$, such that $\|\bm{W}^{*}_{k}-\bm{\hat{X}}_{k}\bm{\hat{X}}_{k}^{\mathsf{H}}\|_{\mathsf{F}}^{2}<\|\bm{W}^{*}_{k}-\bm{X}^{*}_{k}\bm{X}_{k}^{*\mathsf{H}}\|_{\mathsf{F}}^{2}$. Let $\bm{H}_{k}=\sum_{i=1}^{L}h_{k,i}\bm{u}_{k,i}\bm{e}_{i}^{\mathsf{T}}$ and $\bm{\hat{X}}_{k}$ can be expressed as $\bm{\hat{X}}_{k}=\sum_{i=1}^{L}\tilde{s}_{k,i}\bm{u}_{k,i}\bm{e}_{i}^{\mathsf{T}}$, with $\tilde{s}_{k,i}:=s_{k,i}+h_{k,i}$ and $|\tilde{s}_{k,i}|\leq 1$. The difference in the objective value between $\bm{X}^{*}$ and $\bm{\hat{X}}$ is given by:
$\Delta$ $=\|\bm{W}^{*}_{k}-\bm{\hat{X}}_{k}\bm{\hat{X}}_{k}^{\mathsf{H}}\|_{\mathsf{F}}^{2}-\|\bm{W}^{*}_{k}-\bm{X}_{k}^{*}\bm{X}_{k}^{*\mathsf{H}}\|_{\mathsf{F}}^{2}$ (16a)
$=\|\bm{\Sigma}_{k}-{\rm diag}(\tilde{s}_{k,1}^{2},\ldots,\tilde{s}_{k,L}^{2},\ldots,0)\|_{\mathsf{F}}^{2}$
$\quad-\|\bm{\Sigma}_{k}-{\rm diag}(s_{k,1}^{2},\ldots,s_{k,L}^{2},\ldots,0)\|_{\mathsf{F}}^{2}$ (16b)
$=\sum\nolimits_{i=1}^{L}(2\sigma_{k,i}-s_{k,i}^{2}-\tilde{s}_{k,i}^{2})(s_{k,i}^{2}-\tilde{s}_{k,i}^{2})$ (16c)
$\geq\sum\nolimits_{i=1}^{L}(s_{k,i}^{2}-\tilde{s}_{k,i}^{2})^{2}.$ (16d)
Here, the equality in (16b) holds due to the unitary invariance of the Frobenius norm. Noting that $h_{k,i}\neq 0$ for at least one $i$, it follows that $\Delta>0$, leading to a contradiction. Therefore, $\bm{X}_{k}^{*}$ is optimal when $L\leq L_{k}^{*}$.
On the other hand, when $L>L^{*}_{k}$, assume there exists a matrix $\bm{\hat{X}}_{k}=\bm{X}^{*}_{k}+\bm{H}_{k}$ with $\mathrm{rank}(\bm{\hat{X}}_{k})\leq L$ such that $\|\bm{W}^{*}_{k}-\bm{\hat{X}}_{k}\bm{\hat{X}}_{k}^{\mathsf{H}}\|_{\mathsf{F}}^{2}<\|\bm{W}^{*}_{k}-\bm{X}^{*}_{k}\bm{X}_{k}^{*\mathsf{H}}\|_{\mathsf{F}}^{2}$. The matrix $\hat{\bm{X}}_{k}$ can also be expressed as $\bm{\hat{X}}_{k}=\sum_{i=1}^{L}\tilde{s}_{k,i}\bm{u}_{k,i}\bm{e}_{i}^{\mathsf{T}}$. Accordingly, the objective difference can be written as:
$\Delta$ $=\|\bm{W}^{*}_{k}-\bm{\hat{X}}_{k}\bm{\hat{X}}_{k}^{\mathsf{H}}\|_{\mathsf{F}}^{2}-\|\bm{W}^{*}_{k}-\bm{X}_{k}^{*}\bm{X}_{k}^{*\mathsf{H}}\|_{\mathsf{F}}^{2}$ (17a)
$=\|\bm{\Sigma}_{k}-{\rm diag}(\tilde{s}_{k,1}^{2},\ldots,\tilde{s}_{k,L^{*}}^{2},\ldots,\tilde{s}_{k,L}^{2},\ldots,0)\|_{\mathsf{F}}^{2}$
$\quad-\|\bm{\Sigma}_{k}-{\rm diag}(s_{k,1}^{2},\ldots,s_{k,L^{*}_{k}}^{2},\ldots,0)\|_{\mathsf{F}}^{2}$ (17b)
$\geq\sum\nolimits_{i=1}^{L^{*}}(s_{k,i}^{2}-\tilde{s}_{k,i}^{2})^{2}+\sum\nolimits_{i=L^{*}+1}^{L}(\sigma_{k,i}-\tilde{s}_{k,i}^{2})^{2}.$ (17c)
Similarly, $h_{k,i}\neq 0$ for at least one $i$, then $\Delta>0$, again leading to a contradiction. Therefore, $\bm{X}_{k}^{*}$ is also optimal when $L>L_{k}^{*}$. As a result, the solution to $\mathcal{P}_{2}$ is given by:
$\bm{X}^{*}_{k}=\sum\nolimits_{i=1}^{\min(L,L_{k}^{*})}s_{k,i}\bm{u}_{k,i}\bm{e}_{i}^{\mathsf{T}}.$ (18)
This concludes the proof.

### v3

@@ IV-B Power Adaptation of Existing Modulation
Given the signal model in (3), we define $\bm{p}_{\ell}:=[p_{1,\ell},\ldots,p_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ and $\bm{h}_{\ell}:=[h_{1,\ell},\ldots,h_{K,\ell}]^{\mathsf{T}}\in\mathbb{C}^{K}$ as the transmit power vector and channel vector at time slot $\ell$, respectively. Let $\tilde{\bm{x}}_{\ell}$ denotes the modulation vector containing a fixed modulation pattern, i.e., QAM. In the presence of fading, the received constellation point corresponding to the function output $f^{(i)}$ at time slot $\ell$ is then given by:
@@
$\tilde{\bm{v}}^{(i)}_{\ell}=\bm{a}_{i}^{\mathsf{T}}\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})\left(\bm{I}_{K}\otimes\mathds{1}_{Q}\right)\bm{p}_{\ell},$ (14)
@@
where $\otimes$ denotes the Kronecker product. $\bm{I}_{Q}$ and $\bm{I}_{K}$ are identity matrices of size $Q\times Q$ and $K\times K$. $\mathds{1}_{Q}$ is a $Q\times 1$ vector with all elements equal to one, and the operator $\text{diag}(\cdot)$ constructs a diagonal matrix. For simplicity, we let $\bm{B}_{\ell}:=\left(\text{diag}(\bm{h}_{\ell})\otimes\bm{I}_{Q}\right)\text{diag}(\tilde{\bm{x}}_{\ell})(\bm{I}_{K}\otimes\mathds{1}_{Q})$. Accordingly, we formulate the following QCQP to determine the minimum power vectors $\bm{p}_{\ell}$ that satisfiy the constraints in (6):
$\mathcal{P}_{3}:\quad$ $\min_{\{\bm{p}_{\ell}\}\in\mathcal{P}}\quad\sum\nolimits_{\ell=1}^{L}\|\bm{p}_{\ell}\|_{2}^{2}$
${\rm s.t.}\quad$ $\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\geq\Delta f_{ij},\quad\forall(i,j)\in\mathcal{M},$ (15a)
where $\bm{C}_{\ell}^{ij}:=\bm{B}_{\ell}^{\mathsf{H}}(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{B}_{\ell}$ and $\mathcal{P}:=\{\bm{p}_{\ell}\in\mathbb{C}^{K},\ell\in[L]\}$. In a similar manner to $\mathcal{P}_{0}$, we translate the quadratic constraints in $\mathcal{P}_{3}$ into hinge-loss penalties:
@@
$u_{ij}(\{\bm{p}_{\ell}\})=\left[\Delta f_{ij}-\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\right]^{+},\quad\forall(i,j)\in\mathcal{M},$
which measure the constraint violations. At each SCA iteration $s$, we linearize these non-convex terms at the reference points $\{\bm{\tilde{p}}_{\ell}^{(s)}\}$, leading to the following convex subproblem:
$\mathcal{P}_{4}:\min_{\{\bm{p}_{\ell}\}\in\mathcal{P}}\sum\nolimits_{\ell=1}^{L}\bm{p}_{\ell}^{\mathsf{H}}\bm{I}_{K}\bm{p}_{\ell}+\frac{1}{\tilde{M}}\sum_{(i,j)\in\mathcal{M}}w_{ij}^{(s)}(\{\bm{p}_{\ell}\}),$ (16)
where the convex majorizer is defined as:
$w_{ij}^{(s)}(\{\bm{p}_{\ell}\}):=$ $\left[\Delta f_{ij}-\sum\nolimits_{\ell=1}^{L}2\Re\{(\bm{\tilde{p}}_{\ell}^{(s)})^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{p}_{\ell}\}\right$.
$\left.+\sum\nolimits_{\ell=1}^{L}(\bm{\tilde{p}}_{\ell}^{(s)})^{\mathsf{H}}\bm{C}_{\ell}^{ij}\bm{\tilde{p}}_{\ell}^{(s)}\right]^{+}.$ (17)
For the employment of SSD, a subgradient of $w_{ij}^{(s)}(\{\bm{p}_{\ell}\})$ with respect to each variable $\bm{p}_{\ell}$ is given by
$\partial_{\bm{p}_{\ell}}w_{ij}^{(s)}\left(\{\bm{p}_{\ell}\}\right)=\begin{cases}-2\bm{C}_{\ell}^{ij}\bm{\tilde{p}}_{\ell}^{(s)},&\text{if }w_{ij}^{(s)}\left(\{\bm{p}_{\ell}\}\right)>0,\\ \bm{0},&\text{otherwise}.\end{cases}$ (18)
Overall, the complete procedure for solving $\mathcal{P}_{3}$ is summarized in Algorithm 2.
Algorithm 2 SCA-SSD for solving Problem $\mathcal{P}_{3}$
1:Init: SCA iterations $S$, subgradient steps $T$, step sizes $\{\alpha_{t}\}_{t=0}^{T-1}$. Randomly generate initial points $\{\bm{\tilde{p}}_{\ell}^{(0)}\}\in\mathcal{P}$.
2:for $s=0,1,\dots,S$ do
3:for $t=0,1,\dots,T-1$ do
4:  Sample a single pair $(i_{t},j_{t})$ uniformly from $\mathcal{M}$
5:  $\bm{G}_{\ell}^{(t)}=2\bm{p}_{\ell}^{(t)}+\partial_{\bm{p}_{\ell}}w_{ij}^{(s)}(\{\bm{p}_{\ell}^{(t)}\}),\forall\ell\in[L]$
6:  $\bm{p}_{\ell}^{(t+1)}=\bm{p}_{\ell}^{(t)}-\alpha_{t}\,\bm{G}^{(t)}_{\ell},\forall\ell\in[L]$
7:end for
8: Set $\bm{\tilde{p}}_{\ell}^{(s+1)}=\bm{p}_{\ell}^{(T)},\forall\ell\in[L]$
9:end for
10:Output: $\hat{\bm{p}_{\ell}}\leftarrow\bm{\tilde{p}}_{\ell}^{(S+1)},\forall\ell\in[L]$
@@ V Numerical Experiments
where $N_{s}=100$ is the number of Monte Carlo trials, $f^{(i)}$ is the desired function value, $\hat{f}^{(i)}_{j}$ is the estimated value in the $j$-th Monte Carlo trial. $f_{{\rm max}}$ and $f_{{\rm min}}$ denote the maximum and minimum values of the function output, respectively. The fading coefficients $h_{k,\ell}$ are modeled as independent complex Gaussian random variables, i.e., $h_{k,\ell}\sim\mathcal{CN}(0,1)$. For the SSD updates, we adopt a diminishing step-size schedule $\alpha_{t}=\alpha_{0}/\sqrt{t+1}$, where $\alpha_{0}=0.5$.
V-A Performance of SeMAC
For the first experiment, we consider a network of $K=8$ nodes and evaluate the computation of the sum function $f=\sum_{k=1}^{K}x_{k}$ and the product function $f=\prod_{k=1}^{K}x_{k}$, where $x_{k}\in{1,2,3,4}$. The results in Figure 2 show that the NMSE consistently decreases as the Signal-to-Noise Ratio (SNR) increases. In addition, increasing the number of time slots from $L=1$ to $L=2$ improves the computation accuracy by providing greater flexibility in constellation design. Moreover, our proposed SCA–SSD method generates solutions that achieve lower NMSE than the SDR-based method, as it directly refines solutions via minimizing a hinge-loss penalization within the original problem space, thereby avoiding the relaxation errors introduced by lifting and randomization in SDR.
@@ V-A Performance of SeMAC
Figure 2: Performance of SeMAC in terms of NMSE across SNRs using the SDR baseline and the proposed SCA–SSD algorithm.
@@ V-B Comparison to ChannelComp, ReMAC and Bit-Slicing
In this subsection, we compare ChannelComp with repetition, SeMAC, ReMAC, and Bit-Slicing for computing the product function with $K=4$ nodes over $L=2$ time slots, where input values are selected from $x_{k}\in\{1,2,\ldots,16\}$. As shown in Figure 3, by exploiting diverse modulation patterns across time slots, SeMAC achieves lower NMSE by approximately 14 dB compared to ReMAC. While ChannelComp with repetition and ReMAC offer similar performance, ReMAC achieves greater energy efficiency by selectively repeating modulated symbols. Moreover, Bit-Slicing yields the highest NMSE, due to approximation errors introduced during the pre- and post-processing required to support nonlinear functions, as previously analyzed in [6].
@@ V-C Performance of SeMAC-PA
We analyze the performance of SeMAC-PA under Rayleigh fading with $K=8$ and QAM modulation of orders $Q=4$ and $Q=16$ for computing the sum function over $L=2$ time slots. Figure 4 shows that NMSE decreases with increasing SNR, demonstrating the effectiveness of power adaptation. However, SeMAC-PA relies on fixed QAM constellations, making it less effective than the optimized modulation design in SeMAC under channel inversion power control.
@@ VI Conclusion
In this letter, we proposed SeMAC, a multi-symbol modulation framework for digital AirComp that enables reliable function computation through distinct modulation patterns across multiple time slots. To support scenarios with fixed modulation, we introduced SeMAC-PA, which adapts transmit power and phase to reshape the received constellation. One possible future work is to extend SeMAC to scenarios with imperfect CSI and investigate its integration with machine learning techniques for adaptive modulation design.